OpenStudy (anonymous):

Write an equation of a line that is perpendicular to y=-5x+2 and passes through (10,8)

OpenStudy (anonymous):

equation of a line perpendicular to y=mx+c and passing through x1,y1 is given by y-y1= (-1/m) *(x-x1)

OpenStudy (anonymous):

for y=-5x +2 m=-5

OpenStudy (blank ):

y = (x/5)+6 Tell me if you need the steps.

OpenStudy (anonymous):

I would really appreciate the steps @Blank.

OpenStudy (blank ):

Just a second...

OpenStudy (blank ):

y=-5x+2 (10,8) To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept. y=mx+b Using the y=mx+b formula, m=-5. m=-5 The equation of a perpendicular line to y=-5x+2 must have a slope that is the negative reciprocal of the original slope. mperp=-(1)/(-5) The negative reciprocal of <Z>I<z> is 0. mperp=(1)/(5) Find the equation of the perpendicular line using the point-slope formula. (10,8) m=(1)/(5) Find the value of b using the formula for the equation of a line. y=mx+b Substitute the value of m into the equation. y=(1/5)*x+b Substitute the value of x into the equation. y=(1/5)*(10)+b Substitute the value of y into the equation. (8)=(1/5)*(10)+b Since b is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. (1/5)*(10)+b=(8) Multiply ((1)/(5)) by (10) to get ((1)/(5))(10). ((1)/(5))(10)+b=8 Multiply (1)/(5) by 10 to get 2. 2+b=8 Since 2 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2 from both sides. b=-2+8 Add 8 to -2 to get 6. b=6 Now that the values of m(slope) and b(y-intercept) are known, substitute them into y=mx+b to find the equation of the line. y=(x)/(5)+6 Hope this helps. :)

OpenStudy (anonymous):

Thank you!! @Blank 

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)